# A Two-Stage Band-Split Mamba-2 Network for Music Separation

### v1

@@ II-B BMAMBA2-DualNet
As shown in Fig. 2, the BMAMBA2 block contains a forward Mamba2 block and a backward Mamba2 block that model the sequence in forward and backward directions, respectively, and are then concatenated along the N dimension. The forward and backward Mamba-2 blocks are designed as in [28]. Mamba-2 is designed based on the SSD framework, whose core layer is a refinement of Mamba’s selective SSM. For discrete-time signals in digital devices, the discrete SSM updates at each time step $t$ with a hidden state $\mathbf{h}\in\mathbb{R}^{H}$, discretized matrices $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$, as shown in (1) and (2).
$\mathbf{h}_{t}$ $=\overline{\mathbf{A}}\mathbf{h}_{t-1}+\overline{\mathbf{B}}\mathbf{x}_{t},\quad\mathbf{y}_{t}=\mathbf{C}\mathbf{h}_{t}$ (1)
$\overline{\mathbf{A}}$ $=\exp(\Delta\mathbf{A}),\quad\overline{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\Delta\mathbf{B}$ (2)
Here, $\mathbf{A}\in\mathbb{R}^{H\times H}$ is a state transition matrix, $\mathbf{B}\in\mathbb{R}^{H\times 1}$ is an input projection matrix, and $\mathbf{C}\in\mathbb{R}^{1\times H}$ is an output projection matrix, where $H$ is the dimension of the hidden state. $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ are approximated by zero-order hold discretization formulas. A learnable parameter $\Delta$ balances how much to focus on or ignore the current state and input.
@@
The SSD layer is viewed as a map from $\mathbf{A},\mathbf{x},\mathbf{B},\mathbf{C}\mapsto\mathbf{y}$, which produces $\mathbf{A},\mathbf{x},\mathbf{B},\mathbf{C}$ in parallel with a single projection at the beginning of the block. It is worth mentioning the similarity to standard attention architectures in Transformers [26], where $\mathbf{x},\mathbf{B},\mathbf{C}$ correspond to the $\mathbf{Q},\mathbf{K},\mathbf{V}$ projections that are created in parallel.
@@ II-D Fusion Module
As shown in Fig. 1 (E), $D_{1}$ $\in\mathbb{R}^{2\times N\times K\times T}$ is the output from Dual-path Module of the first stage, and $D_{2}$ $\in\mathbb{R}^{2\times N\times K\times T}$ is the output from band split module of the second stage. We concatenate them along the $N$ dimension to form $\hat{D}$ $\in\mathbb{R}^{2\times 2N\times K\times T}$. After passing through a FC layer followed by a Tanh activation function, which reduces the $2N$ dimension back to $N$, $\bar{D}\in\mathbb{R}^{2\times N\times K\times T}$ is obtained.
@@ III-B Model configuration
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{stage1}}+\mathcal{L}_{\text{stage2}}$ (3)
@@
$+|\text{ISTFT}(\mathbf{S}^{i})-\text{ISTFT}(\bar{\mathbf{S}}^{i})|_{1}$ (4)
@@ IV Results and Analysis
Figure 3: Spectrogram examples from the two-stage outputs of the vocal track.

### v2

@@ II-B BMAMBA2-DualNet
As shown in Fig. 2, the BMAMBA2 block contains a forward Mamba2 block and a backward Mamba2 block that model the sequence in forward and backward directions, respectively, and are then concatenated along the N dimension. The forward and backward Mamba-2 blocks are designed as in [28]. Mamba-2 is designed based on the SSD framework, whose core layer is a refinement of Mamba’s selective SSM. Inspired by the continuous system, SSM maps an input function $\mathbf{x}(t)\in\mathbb{R}$ to an output $\mathbf{y}(t)\in\mathbb{R}$ through a higher-dimensional latent state $\mathbf{h}(t)\in\mathbb{R}^{{H}}$ over all continuous time steps $T$, where ${H}$ denotes the number of feature channels. For discrete-time signals in digital devices, the discrete selective SSM updates at each time step $t$ with discretized matrices $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$, as shown in (1) and (2).
$\mathbf{h}_{t}$ $=\overline{\mathbf{A}}\mathbf{h}_{t-1}+\overline{\mathbf{B}}\mathbf{x}_{t},\quad\mathbf{y}_{t}=\mathbf{C}\mathbf{h}_{t}.$ (1)
$\overline{\mathbf{A}}$ $=\exp(\Delta\mathbf{A}),\quad\overline{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\Delta\mathbf{B}.$ (2)
Here, $\mathbf{A}\in\mathbb{R}^{H\times H}$ denotes the evolution parameter, $\mathbf{B}\in\mathbb{R}^{H\times 1}$ and $\mathbf{C}\in\mathbb{R}^{H\times 1}$ denote the projection parameters. $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ are approximated by zero-order hold discretization formulas. A learnable parameter $\Delta$ balances how much to focus on or ignore the current state and input.
@@ II-D Fusion Module
As shown in Fig. 1 (E), $D_{1}$ $\in\mathbb{R}^{2\times N\times K\times T}$ is the output from Dual-path Module of the first stage, and $D_{2}$ $\in\mathbb{R}^{2\times N\times K\times T}$ is the output from band split module of the second stage. We concatenate them along the $N$ dimension to form $\hat{D}$ $\in\mathbb{R}^{2\times 2N\times K\times T}$. After passing through a FC layer followed by a $\tanh$ activation function, which reduces the $2N$ dimension back to $N$, $\bar{D}\in\mathbb{R}^{2\times N\times K\times T}$ is obtained.
@@ III-B Model configuration
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{stage1}}+\mathcal{L}_{\text{stage2}}.$ (3)
@@
$+|\text{ISTFT}(\mathbf{S}^{i})-\text{ISTFT}(\bar{\mathbf{S}}^{i})|_{1},$ (4)
@@ IV Results and Analysis
Figure 3: Spectrogram examples of the vocal track.